A 2.0-kg block slides along a rough horizontal surface. A force(P=6.0N) is applied to the block at a 30degree angle. What is the horizontal acceleration of the block? ASSUME NO FRICTIONAL FORCE
Answer #1

Data Given Mass of the block m = 2.0 kg, Force applied P = 6.0 N at 300

So force will have two components, As shown in the diagram below


P.Sin.30° 300 P.Cos.30o mg

So only horizontal component of force P will cause the motion, So using Newton's second law


F = ma = P\cos 30^{0}

(2kg)a = 6.0 N\cos 30^{0} = 5.196 N

a = \frac{ 5.196 N}{2 kg} = 2.598 m/s^{2} \approx 2.6 m/s^{2}